OpenStudy (anonymous):

3 ( n – 4 ) = 5 ( n + 9 )

OpenStudy (anonymous):

\[3(n-4) = 5(n+9)\]

OpenStudy (anonymous):

Step 2: Use Distributive property \[3n-12 = 5n + 45\]

OpenStudy (anonymous):

Step 3: Combine terms\[2n = -57\]

OpenStudy (anonymous):

Step 4: Solve for n\[n = -28.5\]

OpenStudy (anonymous):

Thank you!!

OpenStudy (denisaboichuk):

−28.5